LunaBechtold:

Guys, Can someone help with this problem??? 10x + 5y = -15

toga:

The equation 10x + 5y = -15 represents a linear equation in two variables x and y. To graph this equation, we can first simplify it by dividing both sides by 5, which gives us 2x + y = -3. This equation represents a line in the coordinate plane, with a slope of -2 and a y-intercept of -3. We can plot two points on this line - one by setting x = 0 and solving for y, and the other by setting y = 0 and solving for x. Connecting these two points with a straight line gives us the graph of the equation.

LunaBechtold:

@toga wrote:
The equation 10x + 5y = -15 represents a linear equation in two variables x and y. To graph this equation, we can first simplify it by dividing both sides by 5, which gives us 2x + y = -3. This equation represents a line in the coordinate plane, with a slope of -2 and a y-intercept of -3. We can plot two points on this line - one by setting x = 0 and solving for y, and the other by setting y = 0 and solving for x. Connecting these two points with a straight line gives us the graph of the equation.
thxxxxxxxxx

toga:

@lunabechtold wrote:
@toga wrote:
The equation 10x + 5y = -15 represents a linear equation in two variables x and y. To graph this equation, we can first simplify it by dividing both sides by 5, which gives us 2x + y = -3. This equation represents a line in the coordinate plane, with a slope of -2 and a y-intercept of -3. We can plot two points on this line - one by setting x = 0 and solving for y, and the other by setting y = 0 and solving for x. Connecting these two points with a straight line gives us the graph of the equation.
thxxxxxxxxx
np

toga:

tag me if you need more help

LunaBechtold:

kk